7.)Solve by completeing the square.
2x^2+x-6=0

Answer by nerdybill(7384)

2x^2+x-6=0
Group terms first:
(2x^2+x)-6=0
(2x^2+x) = 6
2(x^2 + (1/2)x) = 6
(x^2 + (1/2)x) = 3
(x^2 + (1/2)x + 1/16) - 1/16 = 3
(x^2 + (1/2)x + 1/16) = 3 + 1/16
(x^2 + (1/2)x + 1/16) = 48/16 + 1/16
(x + 1/4)^2 = 49/16
Taking the square root of both sides:
x + 1/4 = (+-)7/4
x = -1/4 + (+-)7/4
.
Two possible solutions then are:
x = -1/4 + 7/4 = 6/4 = 1.5
or
x = -1/4 - 7/4 = -8/4 = -2